# Isometric rigidity of the Wasserstein space over the plane with the maximum metric

### v1

@@ Abstract.
We study $p$-Wasserstein spaces over the branching spaces $\mathbb{R}^{2}$ and $[0,1]^{2}$ equipped with the maximum norm metric. We show that these spaces are isometrically rigid for all $p\geq 1,$ meaning that all isometries of these spaces are induced by isometries of the underlying space via the push-forward operation. This is in contrast to the case of the Euclidean metric since with that distance the $2$-Wasserstein space over $\mathbb{R}^{2}$ is not rigid. Also, we highlight that $1$-Wasserstein space is not rigid over the interval $[0,1]$, while according to our result, its two-dimensional analog with the more complicated geodesic structure is rigid. To the best of our knowledge, our results are the first ones concerning the structure of isometries of Wasserstein spaces over branching spaces. Accordingly, our proof techniques substantially deviate from those used for non-branching underlying spaces.
@@
Z. M. Balogh is supported by the Swiss National Science Foundation, Grant Nr. 200020_191978.
T. Titkos is supported by the Hungarian National Research, Development and Innovation Office - NKFIH grant no. K134944) and by the Momentum program of the Hungarian Academy of Sciences under grant agreement no. LP2021-15/2021.
D. Virosztek is supported by the Momentum program of the Hungarian Academy of Sciences under grant agreement no. LP2021-15/2021, and partially supported by the ERC Synergy Grant No. 810115.
G. Kiss is supported by János Bolyai Research Fellowship of the Hungarian Academy of Sciences and by the Hungarian National Research, Development and Innovation Office - NKFIH grants no. K146922, FK142993.
@@ 1. Introduction and the main result
In this paper, we consider the situation of branching spaces, namely $\mathbb{R}^{2}$ and $Q=[0,1]^{2}$ endowed with the maximum metric. Since the above-mentioned technique does not work in our case, we shall use a different method in order to determine the structure of the isometry group of the Wasserstein space over these spaces.
@@
where $C(\mu,\nu)$ is in the set of couplings between $\mu$ and $\nu$. That is, $\pi\in\mathcal{P}(X\times X)$ and its marginals are equal to $\mu$ and $\nu$: $\pi(A\times X)=\mu(A)$, and $\pi(X\times A)=\nu(A)$ for any Borel set $A\subseteq X$. For the sake of brevity, we will denote the Wasserstein space $\big{(}\mathcal{P}_{p}(X,d_{\mathrm{{m}}}),d_{W_{p}}\big{)}$ by $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$. The support of a measure $\mu$ will be denoted by $\mathrm{supp}(\mu)$. For some distinguished collection of lines $L\subset\mathbb{R}^{2}$ the set
$\mathcal{W}_{p}(L,d_{\mathrm{{m}}})=\{\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}}):\mathrm{supp}(\mu)\subseteq L\}$
@@
Let us recall that by the results of Kloeckner [12] the quadratic Wasserstein space $\mathcal{W}_{2}(\mathbb{R}^{n},d_{\|\cdot\|_{2}})$ is not rigid as it has non-trivial shape-preserving isometries. Moreover, in the case $n=1$ there is a flow of exotic (non-shape-preserving) isometries. Furthermore, the structure of the isometry group of Wasserstein spaces could depend both on the choice of $X$ and the value of $p$. Indeed, the results of [9] show in the one-dimensional case $X=\mathbb{R}$ that the isometry group of $\mathcal{W}_{2}(\mathbb{R},d_{|\cdot|})$ is much larger than the isometry group of $\mathcal{W}_{p}(\mathbb{R},d_{|\cdot|})$ for all $p\neq 2$, while if $X=[0,1]$, then the isometry group of $\mathcal{W}_{1}([0,1],d_{|\cdot|})$ is richer than the isometry group of $\mathcal{W}_{p}([0,1],d_{|\cdot|})$ for all $p>1$ (see [9]).
In this paper, we distinguish the cases $p=1$ and $p>1$. We note that the case $p<1$ has already been covered by the general result [10, Corollary 4.7.] which says that the Wasserstein space $\mathcal{W}_{p}(X,d)$ is isometrically rigid by for every Polish underlying space $(X,d)$ and for every parameter $p<1.$ Furthermore, the underlying space $X$ will be either $\mathbb{R}^{2}$ or the unit square $Q=[0,1]^{2}$. Our main result shows, that in contrast to the above non-rigidity results in the one-dimensional case, (and also in the higher dimensional $\mathbb{R}^{n}$ with the Euclidean metric) in our situation the Wasserstein spaces are isometrically rigid when the underlying space $\mathbb{R}^{2}$ or $Q$ is considered with the maximum metric.
@@
Let $X=\mathbb{R}^{2}$ or $X=Q=[0,1]^{2}$ equipped with the maximum metric. Then for any $p\geq 1$ the Wasserstein space $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is isometrically rigid. That is, for any isometry $\Phi:\mathcal{W}_{p}(X,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ there exists a unique isometry $T:(X,d_{\mathrm{{m}}})\to(X,d_{\mathrm{{m}}})$ such that
@@
$L_{+}=\{(t,t)\colon t\in\mathbb{R}\}\quad\text{and}\quad L_{-}=\{(t,-t):t\in\mathbb{R}\},$
@@
We call the Wasserstein space $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ diagonally rigid, if all of its isometries preserve diagonally supported measures up to composition by a trivial isometry. More precisely, for every isometry $\Phi:\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\to\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ there exists an isometry $T:(\mathbb{R}^{2},d_{\mathrm{{m}}})\to(\mathbb{R}^{2},d_{\mathrm{{m}}})$ such that $\Phi(\mu)=T_{\#}(\mu)$ whenever the support of $\mu$ is a subset of $L_{+}$ or $L_{-}$.
The concept of diagonal rigidity can also be introduced for the space $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ by considering translated copies of $L_{+}$ and $L_{-}$ going through the point $(\frac{1}{2},\frac{1}{2})$ as the appropriate diagonals for $Q$.
Our aim in Section 2 is to prove that diagonal rigidity implies rigidity. This step is quite general in the sense that its proof works the same way for any $p\geq 1$ and $X=\mathbb{R}^{2}$ or $X=[0,1]^{2}$. To prove that $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is indeed diagonally rigid is more tricky and uses very different arguments for different underlying spaces $X=\mathbb{R}^{2}$ and $X=[0,1]^{2}$ and for different values of $p$. These results are proven in Section 3.
@@ 2. Diagonal rigidity implies rigidity
Our first result shows that isometries of the Wasserstein space $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ are diagonally determined in the sense that if an isometry acts on diagonally supported measures as a trivial isometry, then it is a trivial isometry. More precisely, if $\Phi$ is an isometry such that $\Phi(\mu)=T_{\#}(\mu)$ for every diagonally supported measure, then $\Phi=T_{\#}$. In this part, it does not make a difference to consider the space $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ or $\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ and the value of $p\geq 1$ is also arbitrary.
@@
Let $p\geq 1$ and let $\Phi:\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\to\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ be an isometry, such that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(L_{+},d_{\mathrm{{m}}})\cup\mathcal{W}_{p}(L_{-},d_{\mathrm{{m}}})$. Then $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}}).$ And similarly, if $\Phi:\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ is an isometry, such that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(L_{1,0}\cap Q,d_{\mathrm{{m}}})\cup\mathcal{W}_{p}(L_{-1,1}\cap Q,d_{\mathrm{{m}}})$. Then $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(Q,d_{\mathrm{{m}}}).$
For the sake of brevity, we only prove the case $X=\mathbb{R}^{2}$. The same argument works in the case $X=Q$, replacing lines $L$ by line segments $L\cap Q$ and by replacing $L_{+}$ and $L_{-}$ by the diagonals of the square, i.e., $L_{1,0}\cap Q$ and $L_{-1,1}\cap Q$, respectively. The proof of the statement will be a combination of lemmas. The first lemma is about the minimal distance projection onto lines.
@@
If $x\in L$, then we take $\hat{x}:=x$ and the claim is obvious. If $x\not\in L$, then there is a small enough $r>0$ such that $B(x,r)\cap L=\emptyset$, where by $B(x,r)$ we denote the open metric ball with center $x$ and radius $r>0$ with respect to the maximum metric $d_{\mathrm{{m}}}$.
Now we define $R:=\sup\{r>0:B(x,r)\cap L=\emptyset\}$. Then $\bar{B}(x,R)\cap L\neq\emptyset$, moreover, $\bar{B}(x,R)\cap L\subseteq\partial B(x,R)$. Furthermore, since both $\bar{B}(x,R)$ and $L$ are convex sets, it follows that their intersection $\bar{B}(x,R)\cap L$ is also convex, and thus it consists of either one single point or a line segment.
Since this intersection is a subset of the boundary of $\bar{B}(x,R)$, thus it can be a line segment only if $L$ is parallel to the $x$-axis or the $y$-axis. By assumption, this is not the case. Therefore $\bar{B}(x,R)\cap L$ is a single point. Denote $\hat{x}=\bar{B}(x,R)\cap L$. It is easy to check that $\hat{x}$ is the unique point in $L$ that satisfies the condition $d_{\mathrm{{m}}}(x,\hat{x})\leq d_{\mathrm{{m}}}(x,y)$ for all $y\in L.$
@@
Let $\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$. Then, $\hat{\mu}={P_{L}}_{\#}(\mu)$ is the unique measure in $\mathcal{W}_{p}(L,d_{\mathrm{{m}}})$ such that
@@
We prove first that the set
$S=\{(x,P_{L}(x))\colon x\in\mathbb{R}^{2}\}\subseteq\mathbb{R}^{2}\times L$
is cyclically monotone. To see this, let $N\geq 1$ and $(x_{i},P_{L}(x_{i}))_{i=1}^{N}$ be a finite set of points in $S$. We have to check that
(2) $\sum_{i=1}^{N}d_{\mathrm{{m}}}^{p}(x_{i},P_{L}(x_{i}))\leq\sum_{i=1}^{N}d_{\mathrm{{m}}}^{p}(x_{i+1},P_{L}(x_{i})),.$
where $x_{N+1}=x_{1}$. By the definition of $P_{L}$, we have that
(3) $d_{\mathrm{{m}}}^{p}(x_{i},P_{L}(x_{i}))\leq d_{\mathrm{{m}}}^{p}(x_{i},P_{L}(x_{i-1})),\ \text{ for all}\ i=1,\ldots,N$
where $x_{0}=x_{N}$. Taking the sum in the inequalities (3) for $i=1$ to $N$ and relabeling the indices on the right side, implies inequality (2).
Now we apply Corollary 2.6.8 of Figalli-Glaudo [8], which implies that $P_{L}:\mathbb{R}^{2}\to L$ is an optimal transport map between $\mu$ and $\hat{\mu}={P_{L}}_{\#}(\mu)$. Thus
$d_{W_{p}}^{p}(\mu,\hat{\mu})=\int\limits_{\mathbb{R}^{2}}d_{\mathrm{{m}}}^{p}(x,P_{L}(x))~{}\mathrm{d}\mu(x).$
@@
$\begin{split}d_{W_{p}}^{p}(\mu,\nu)&=\int\limits_{\mathbb{R}^{2}\times L}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi(x,y)\\ &\geq\int\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}d_{\mathrm{{m}}}^{p}(x,P_{L}(x))~{}\mathrm{d}\pi(x,y)\\ &=\int\limits_{\mathbb{R}^{2}}d_{\mathrm{{m}}}^{p}(x,P_{L}(x))~{}\mathrm{d}\mu(x)=d_{W_{p}}^{p}(\mu,\hat{\mu}).\end{split}$
@@
To show that $\hat{\mu}$ is the unique minimizer, assume that $\nu\in\mathcal{W}_{p}(L,d_{\mathrm{{m}}})$ is another minimizer. Then it must be equality throughout the above chain of inequalites. This implies that for $\pi$ a.e. $(x,y)$ we must have $(x,y)=(x,P_{L}(x))$ and thus $\pi=(Id\times P_{L})_{\#}(\mu)$, which gives that $\nu={P_{L}}_{\#}\mu=\hat{\mu}$, as required. ∎
@@
After this preparation we can turn to the proof of Proposition 2. The proof is inspired by Bertrand and Kloeckner [5] and it is based on the method of Radon transform. In our case the Radon transform will be a mapping $\mathcal{R}:\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\times\{L_{+},L_{-}\}\to\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ defined by
(4) $\mathcal{R}(\mu,L_{+}):={P_{L_{+}}}_{\#}(\mu),\qquad\qquad\mathcal{R}(\mu,L_{-}):={P_{L_{-}}}_{\#}(\mu).$
Proof of Proposition 2: Assume that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(L_{+},d_{\mathrm{{m}}})\cup\mathcal{W}_{p}(L_{-},d_{\mathrm{{m}}})$. Now we want to extend this to the whole $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$. The main idea is to consider a subset $\mathcal{F}\subset\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ such that

### v3

@@ Abstract.
We study $p$-Wasserstein spaces over the branching spaces $\mathbb{R}^{2}$ and $[-1,1]^{2}$ equipped with the maximum norm metric. We show that these spaces are isometrically rigid for all $p\geq 1,$ meaning that all isometries of these spaces are induced by isometries of the underlying space via the push-forward operation. This is in contrast to the case of the Euclidean metric since with that distance the $2$-Wasserstein space over $\mathbb{R}^{2}$ is not rigid. Also, we highlight that the $1$-Wasserstein space is not rigid over the closed interval $[-1,1]$, while according to our result, its two-dimensional analog, the closed unit ball $[-1,1]^{2}$ with the more complicated geodesic structure is rigid.
@@
Z. M. Balogh is supported by the Swiss National Science Foundation, Grant Nr. 200021_228012. G. Kiss is supported by János Bolyai Research Fellowship of the Hungarian Academy of Sciences and by the Hungarian National Research, Development and Innovation Office - NKFIH grants no. K146922, FK142993, Starting 150576. T. Titkos is supported by the Hungarian National Research, Development and Innovation Office - NKFIH grant no. K134944), by the Momentum program of the Hungarian Academy of Sciences under grant agreement no. LP2021-15/2021., and by the C-PAP Scholarship of Corvinus University. D. Virosztek is supported by the Momentum program of the Hungarian Academy of Sciences under grant agreement nr. LP2021-15/2021, by the Hungarian National Research, Development and Innovation Office (NKFIH) under grant agreement nr. Excellence_151232, and partially supported by the ERC Synergy Grant No. 810115.
@@ 1. Introduction and the main result
In this paper, we consider the situation of branching spaces, namely $\mathbb{R}^{2}$ and $Q=[-1,1]^{2}$ endowed with the maximum metric. Since the above-mentioned technique does not work in our case, we shall use a different method in order to determine the structure of the isometry group of the Wasserstein space over these spaces.
@@
where $C(\mu,\nu)$ is in the set of couplings between $\mu$ and $\nu$. That is, $\pi\in\mathcal{P}(X\times X)$ and its marginals are equal to $\mu$ and $\nu$: $\pi(A\times X)=\mu(A)$, and $\pi(X\times A)=\nu(A)$ for any Borel set $A\subseteq X$. Recall that if $10<p<$ 1, then the definition of the $p$-Wasserstein distance is slightly different. In that case, $d_{W_{p}}(\mu,\nu)=\min_{\pi\in C(\mu,\nu)}\iint\limits_{X\times X}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi(x,y)$.
For the sake of brevity, we will denote the Wasserstein space $\big{(}\mathcal{P}_{p}(X,d_{\mathrm{{m}}}),d_{W_{p}}\big{)}$ by $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$.
The support of a measure $\mu$ will be denoted by $\mathrm{supp}(\mu)$. For some distinguished collection of lines $L\subset\mathbb{R}^{2}$ the set
$\mathcal{W}_{p}(L,d_{\mathrm{{m}}})=\{\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\,|\,\mathrm{supp}(\mu)\subseteq L\}$
@@
Let us recall that by the results of Kloeckner [13] the quadratic Wasserstein space $\mathcal{W}_{2}(\mathbb{R}^{n},d_{\|\cdot\|_{2}})$ is not rigid as it has non-trivial shape-preserving isometries. Moreover, in the case $n=1$ there is a flow of exotic (non-shape-preserving) isometries. Furthermore, the structure of the isometry group of Wasserstein spaces could depend both on the choice of $X$ and the value of $p$. Indeed, the results of [10] show in the one-dimensional case $X=\mathbb{R}$ that the isometry group of $\mathcal{W}_{2}(\mathbb{R},d_{|\cdot|})$ is much larger than the isometry group of $\mathcal{W}_{p}(\mathbb{R},d_{|\cdot|})$ for all $p\neq 2$, while if $X=[0,1]$, then the isometry group of $\mathcal{W}_{1}([0,1],d_{|\cdot|})$ is richer than the isometry group of $\mathcal{W}_{p}([0,1],d_{|\cdot|})$ for all $p>1$ (see [10]). As it was already pointed out in [10], the same conclusion holds for every compact interval $[a,b]$. For our considerations, the relevant conclusion is that $\mathcal{W}_{p}([-1,1],|\cdot|)$ is rigid if and only if $p\neq 1$.
In this paper, we distinguish the cases $p=1$ and $p>1$. We note that the case $p<1$ has already been covered by the general result [11, Corollary 4.7.] which says that the Wasserstein space $\mathcal{W}_{p}(X,d)$ is isometrically rigid for every Polish underlying space $(X,d)$ and for every parameter $p<1.$ Furthermore, the underlying space $X$ will be either $\mathbb{R}^{2}$ or the closed unit ball $Q=[-1,1]^{2}$. Our main result shows, that in contrast to the above non-rigidity results in the one-dimensional case, (and also in the higher dimensional $\mathbb{R}^{n}$ with the Euclidean metric) in our situation the Wasserstein spaces are isometrically rigid when the underlying space $\mathbb{R}^{2}$ or $Q$ is considered with the maximum metric.
@@
Let $X=\mathbb{R}^{2}$ or $X=Q=[-1,1]^{2}$ equipped with the maximum metric. Then for any $p\geq 1$ the Wasserstein space $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is isometrically rigid. That is, for any isometry $\Phi:\mathcal{W}_{p}(X,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ there exists a unique isometry $T:(X,d_{\mathrm{{m}}})\to(X,d_{\mathrm{{m}}})$ such that
@@
$L_{+}=\{(t,t)\,|\,t\in\mathbb{R}\}\quad\text{and}\quad L_{-}=\{(t,-t)\,|\,t\in\mathbb{R}\},$
@@
When we are working in $Q=[-1,1]^{2}$, (with a slight abuse of notation) these symbols denote the line segment contained in $Q$.
@@
Definition 2.
Let $X$ be either $\mathbb{R}^{2}$ or $Q$ equipped by the maximum metric. We call the Wasserstein space $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ diagonally rigid, if for every Wasserstein isometry $\Phi:\mathcal{W}_{p}(X,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ there exists an isometry $T:(X,d_{\mathrm{{m}}})\to(X,d_{\mathrm{{m}}})$ such that $\Phi(\mu)=T_{\#}(\mu)$ whenever the support of $\mu$ is a subset of $L_{+}$ or $L_{-}$.
Our aim in Section 2 is to prove that diagonal rigidity implies rigidity. This step is quite general in the sense that its proof works the same way for any $p\geq 1$ and $X\in\{\mathbb{R}^{2},Q\}$. To prove that $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is indeed diagonally rigid is more tricky and uses very different arguments for different underlying spaces $X=\mathbb{R}^{2}$ and $X=[-1,1]^{2}$ and for different values of $p$. These results are proven in Section 3.
@@ 2. Diagonal rigidity implies rigidity
The main result of this section is the following:
@@
Let $p\geq 1$ and $X\in\{\mathbb{R}^{2},Q\}$. Assume that the space $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is diagonally rigid. Then the $\mathcal{W}_{p}(X,d_{\mathrm{{m}}})$ is rigid.
For the sake of brevity, we only prove the case $X=\mathbb{R}^{2}$. The same argument works in the case $X=Q$, replacing lines by line segments contained in $Q$.
The proof of the statement will be a combination of lemmas. The first lemma is about the minimal distance projection onto lines.
@@
If $x\in L$, then we take $\hat{x}:=x$ and the claim is obvious. If $x\not\in L$, then $\hat{x}$ is the first point of contact of metric balls centered at $x$ with $L$. Since metric balls are squares aligned with the $x$ and $y$-axes, and by assumption $L$ is not parallel to any of these axes, $\hat{x}$ is uniquely defined.
@@
Let $\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$. Then, the measure $\hat{\mu}={P_{L}}_{\#}(\mu)$ is the metric projection $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\to\mathcal{W}_{p}(L,d_{\mathrm{{m}}})$ i.e. the unique measure in $\mathcal{W}_{p}(L,d_{\mathrm{{m}}})$ such that
@@
$\begin{split}d_{W_{p}}^{p}(\mu,\nu)&=\int\limits_{\mathbb{R}^{2}\times L}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi(x,y)\geq\int\limits_{\mathbb{R}^{2}\times\mathbb{R}^{2}}d_{\mathrm{{m}}}^{p}(x,P_{L}(x))~{}\mathrm{d}\pi(x,y)\\ &=\int\limits_{\mathbb{R}^{2}}d_{\mathrm{{m}}}^{p}(x,P_{L}(x))~{}\mathrm{d}\mu(x)=d_{W_{p}}^{p}(\mu,\hat{\mu}).\end{split}$
@@
To show that $\hat{\mu}$ is the unique minimizer, note that in the case equality we have that $y=P_{L}(x)$ for $\pi$ almost every $(x,y)$ showing that $\pi=(Id\times P_{L})_{\#}\mu$ and thus $\nu=P_{L\#}\mu$. ∎
@@
After this preparation we can turn to the proof of Proposition 3. The proof is inspired by Bertrand and Kloeckner [6] and it is based on the method of Radon transform. In our case the Radon transform will be a mapping $\mathcal{R}:\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})\to\mathcal{W}_{p}(L_{+},d_{\mathrm{{m}}})\times\mathcal{W}_{p}(L_{-},d_{\mathrm{{m}}})$ defined by
(2) $\mathcal{R}(\mu):=\Big{(}{P_{L_{+}}}_{\#}(\mu),{P_{L_{-}}}_{\#}(\mu)\Big{)}.$
Proof of Proposition 3.
Without loss of generality we can assume that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(L_{+},d_{\mathrm{{m}}})\cup\mathcal{W}_{p}(L_{-},d_{\mathrm{{m}}})$. Now we want to extend this property to the whole $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$. The main idea is to consider a subset $\mathcal{F}\subset\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ such that